OpenStudy (anonymous):

Model the function rule y = -2x - 3 with a table of values.

OpenStudy (luigi0210):

How many values?

OpenStudy (luigi0210):

It's easier to use:  Now plug that in for x and solve for y

OpenStudy (anonymous):

I don't know.it doesn't say.

OpenStudy (anonymous):

1 -1 -3 -5 -7

OpenStudy (anonymous):

how do you do it?

OpenStudy (luigi0210):

@cutepochacco NO

OpenStudy (luigi0210):

Let the asker learn~

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

Break it down

OpenStudy (luigi0210):

Just plug in the x values and solve, so for -2: \[\LARGE y=-2(-2)-3\] \[\LARGE y=4-3\] \[\LARGE y=1\] And that's what you put in the table:  For the next plug in -1, what do you get? \[\LARGE y=-2(-1)-3\]